# Large Language Models Think Too Fast To Explore Effectively

### v1

@@ 4 Discussion and Conclusion
Paper Summary. Exploration is essential for discovering new opportunities and understanding complex environments. Our study reveals that most LLMs struggle to achieve human-level exploration in the open-ended task. They heavily rely on uncertainty-driven strategies, which provide short-term gains but fail to support long-term success. While both uncertainty and empowerment are present in their latent spaces, LLMs fail to balance these strategies effectively, resulting in suboptimal performance and limited adaptability to broader decision spaces. However, we also find exceptional models like o1, which surpass humans in performance and indicate stronger uncertainty-driven and empowerment exploration strategy usage. This suggests that LLMs with reasoning training may be essential to perform in open-ended tasks, which requires a variety of exploration strategies.
@@
Fast Thinking in Traditional LLMs. A key issue lies in LLMs “thinking too fast” during exploratory tasks. In LLaMA3.1-70B, uncertainty values dominate early transformer blocks and the activations from early transformer blocks correlate strongly with immediate choices, while empowerment values emerge in middle blocks. This temporal mismatch leads to premature decision-making that prioritizes short-term utility over deeper exploration. This predominant information processing by uncertainty values and choices can weaken the role of empowerment in the exploratory decision-making process. We also replicate a similar finding in LLaMA3.1-8B (Figure.13 in the appendix).
We believe that the traditional inference paradigms mainly drive this premature information processing in traditional LLMs. The autoregressive way of generating tokens based on conditional probabilities of existing context may diminish the potential exploration capacity of LLMs. We additionally tried several common ways to improve the model performance, such as prompt engineering, intervention (feature steering), and seeking alternative models (appendix.C). The additional results show that both prompt engineering and intervention did not help to improve the model performance. In the model alternative, we experiment on a most recently released open-source reasoning model, DeepSeek-R1 [5], which claims to match o1-level performance in multiple benchmarks, with its reasoning process visible. This reasoning model outperforms other traditional LLMs and reaches human-level performance in this task (Figure.9 in the appendix). This superior performance provides more evidence that model infrastructure may be the main reason that limits their performance in this type of task.
@@
Lmitations and Future Directions. Despite these findings, the underlying cause of LLMs “thinking too fast” remains unclear and requires further investigation. Future research could explore the interaction between model architecture and processing dynamics, as well as how LLMs weigh uncertainty and empowerment during decision-making. Interventions such as integrating extended reasoning frameworks like CoT, optimizing transformer block interactions, or training with explicit exploratory objectives could enhance LLMs’ exploratory abilities. These efforts would not only improve performance but also advance our understanding of creating AI systems capable of more human-like exploration.
Acknowledgments
We thank Franziska Brändle for helping make up the source data and codes for the human empowerment study. We also thank Can Demircan and Huadong Xiong for the helpful discussion in training and interpreting Sparse AutoEncoders in the paper. This work was funded by a SCIALOG Award #29079 from the Research Corporation for Scientific Advancement (to RCW) and the OpenAI Researcher Access Program (to HX). This research was also supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA. Codes and data will not be publicly available until the paper is peer-reviewed. However, they can be available upon special requests to the authors.
@@ Appendix B The LLM Behavior Across Different Temperatures
To better understand the exact changing behaviors of LLMs under different temperatures, we categorized all combinations into four types: whether this generates a new element, and whether these two combinations have been used before. Our result shows that for even larger models, as temperature increases, the percentage of choosing existing failed combinations decreases. This reveals a diminishing number of redundant behaviors among LLMs. In the meantime, the percentage of choosing new but failed combinations increases significantly, which suggests that the model tends to choose new combinations more often in higher temperatures than in lower temperatures. More importantly, a majority of these new combinations do not generate new elements, suggesting that the high temperatures only alter the uncertainty-driven exploration strategies but not empowerment, since in larger spaces, only random combinations are not sufficient to perform the task effectively. This also explains why higher temperatures can moderately improve the model performance but are still distant from human behaviors.
@@ Appendix C Attempts for Model Improvements
To investigate any potential general way to improve the performance of models, we had several attempts including prompt engineering, interventions, and experiments on alternative open-source models.
@@ C.1 Prompt Engineering
Because the model exhibited repeated behaviors and did not fully utilize “empowerment”, we introduced more direct, guiding prompts (including the steps highlighted in Figure 7) to help gpt-4o($temperature=1$) make more diverse and forward-looking choices, including Chain-of-Thought (CoT, Wei et al. [18]), self-reflection [20] and even explicit strategy hint. However, the result shows that simply updating the prompt—even with explicit instructions and reasoning strategies—didn’t have a significant improvement in performance($average\ inventory=43,t=-1.17,p=0.304$). The model continued to repeat combinations and did not significantly increase its ability to discover new elements. It averages 43 elements. This suggests that while prompt engineering can help shape a model’s outputs, it may not be enough on its own to overcome certain ingrained tendencies such as repeating prior actions or failing to maximize exploratory behavior.
@@ C.2 Open-Source Reasoning Model - DeepSeek-R1
The result shows that DeepSeek-R1 reached near human-level task performance (Figure.9A), but still underperforms than o1. In the behavioral patterns (Figure.9B), DeepSeek-R1, compared to traditional LLMs have fewer attempts on existing combinations, showing a stronger exploration strategy usage. However, compared to humans and o1, DeepSeek-R1 tried more on failed new element combinations but not successful ones, suggesting DeepSeek-R1 may explore less effectively than humans and o1. Our regression results confirm furtherly that DeepSeek-R1 exhibit stronger/weaker uncertainty-driven exploration stratgies and stronger/weaker empowerment, echoing its underperformance than o1.
Figure 9: A: Best Temperature of Each Model and Human Performance. B: Best Temperature of Each Model and Human Behaviors. Choose the LLM models’(gpt-4o, LLaMA3.1-8B, LLaMA3.1-70B) best performance at temperature = 1, and compare it with human and o1, deepseek-reasoner(temperature = 1). Compare each model’s performance and behaviors.C: Regression estimates by temperature and model.
To conduct a deeper investigation, we also collected the model’s reasoning process along with the experiment. Here, we use a qualitative analysis of some pieces of its reasoning process. For example, in trial 200, the model discovered 48 elements. The specific reasoning process is in Figure. 10.
The model engages in a systematic multi-stage reasoning process to propose a new combination from its inventory. It begins by reviewing the inventory and cross-referencing prior failed attempts to avoid redundancy. The model revisits successful paths (e.g., how ’village’ was created from ’house + field’) and evaluates logical relationships(empowerment) between elements, such as exploring combinations involving larger structures like ’continent.’ But the model doesn’t choose a combination containing ’continent’. It discards irrelevant or infeasible paths, such as those requiring missing elements (’ash,’ ’fruit’). After confirming the presence of both ’city’ and ’water’ in the inventory, it identifies that ’city + water’ has not been tested and hypothesizes that this combination could yield a new element, such as ’port’ or ’harbor.’ This iterative and logic-driven process highlights the model’s ability to balance memory, deduction, and validation in problem-solving. However, the shortcoming is obvious. Although there are many logical expressions, the model’s reasoning process is not straightforward to its final choice, where we only see the last sentence in red can somehow explain the underlying motivation. This could probably suggest that the model’s reasoning process is somehow redundant and may needs further refinement to make efficient decisions.
@@
Figure 10: The reasoning process of DeepSeek-R1 in Trial 200: The model explores possible combinations(blue color part) for discovering a new element in the game by systematically reviewing inventory elements(orange color part), prior attempts(blue color part), and logical inferences to make the decision(red color part). It also found some more empowerment elements(green color part) but didn’t choose them.
@@
To investigate the role of empowerment and uncertainty representations in LLaMA3.1-70B during in-context learning, we analyzed the most correlated neurons with these values and performed targeted interventions to assess their impact on model performance. Given that both empowerment and uncertainty were highly demanding in the task (shown by the example of o1’s superior performance), we are wondering whether intervention to strengthen both representations of uncertainty and empowerment could generate better performance. We found that interventions in the uncertainty layer, which is located relatively early in the model (similar to the choice layer), caused severe performance degradation, with even minor adjustments rendering LLaMA3.1-70B unable to perform the task effectively. This suggests that uncertainty representation plays a fundamental role in processing task context and history in in-context learning. Conversely, enhancing the empowerment layer also can’t improve the performance. Even when set to an intervention factor of 1.5, which brought the model closer to the original level, other intervention values resulted in decreased performance. Additionally, ablation experiments, in which the most correlated neuron activations were zeroed out, further confirmed the critical role of these representations. Therefore, as highlighted before, we argue the limit in infrastructure in traditional LLMs is the main reason for their below-human-level performance in open-ended exploration tasks.
Figure 11: LLaMA3.1-70B Average Inventory of Uncertainty Intervention. Set 5 different levels of uncertainty intervention(0.0, 0.5, 0.7, 1.0, 2.0). Increasing the uncertainty intervention progressively disrupts the model’s ability to complete the task, indicating the critical role of early uncertainty layers in processing task history and context. LLaMA3.1-70B Average Inventory of Empowerment Intervention. Performance remains closer to the original level when the intervention is set to 1.5, whereas other levels of intervention result in performance degradation.
@@ Appendix D SAE Setup
We investigated the role of the empowerment and uncertainty layers in LLaMA3.1-8B (temperature = 1) by training a Sparse Autoencoder (SAE) to identify and interpret their activations, followed by targeted interventions where each layer’s activation was set to zero. The results, summarized in Figure.13A and Figure.13B, show that setting the empowerment layer to zero had a minimal effect on regression estimates and only slightly reduced model performance, suggesting that the empowerment layer has a limited role in sustaining task performance. In contrast, setting the uncertainty layer to zero led to a substantial reduction in regression estimates for uncertainty, accompanied by a marked decline in model performance. This highlights the critical importance of the uncertainty layer in facilitating exploration and maintaining robust decision-making capabilities within the model.
Figure 13: A: SAE Correlation Analysis. Maximum correlation of uncertainty values across layers, peaking at layer 7. Maximum correlation of empowerment values across layers, peaking at layer 22. Maximum correlation of choices across layers, peaking at layer 15. B: LLaMA3.1-8B Intervention Regression Results.The bars represent the regression estimates for empowerment, and uncertainty under the original condition(LLaMA3.1-8B, temperature = 1), empowerment intervention (set to zero), and uncertainty intervention (set to zero). C: LLaMA3.1-8B Average Inventory of Interventions. Uncertainty intervention leads to a significant reduction in the average inventory, indicating its essential role in model performance.

### v2

@@ 5 Discussion and Conclusion
Paper Summary. Exploration is critical for navigating complex, open-ended tasks, yet most LLMs fall short of human-level performance. They tend to over-rely on uncertainty-driven strategies that yield short-term gains but hinder long-term success. Although both uncertainty and empowerment signals are present in their latent spaces, LLMs generally fail to balance them effectively. Notably, models like o1 outperform humans, suggesting that reasoning-trained LLMs can better leverage diverse exploration strategies. Our thought analysis further shows that reasoning models engage in deeper, more varied reasoning and allocate more tokens to deliberation, indicating more effortful decision-making than standard LLMs.
@@
Fast Thinking in Traditional LLMs. A key limitation of traditional LLMs in exploratory tasks is their tendency to “think too fast.” In LLaMA3.1-70B, we observe that uncertainty-related signals dominate the early transformer layers and correlate strongly with immediate choices, while empowerment signals emerge only in the middle layers. This temporal mismatch causes premature decision-making, favoring short-term utility over sustained exploration. A similar pattern is replicated in LLaMA3.1-8B (Appendix Figure 16).
We argue that this limitation is largely driven by the autoregressive inference paradigm itself, where next-token generation is conditioned on shallow context accumulation. This architecture may inherently suppress deeper or delayed exploration. We evaluated several mitigation strategies, including prompt engineering and activation intervention (feature steering) (Appendix C). Neither prompt tuning nor intervention significantly improved performance.
In contrast, the reasoning model DeepSeek-R1 [9] and o1 [22], which explicitly support multi-step reasoning, reached human-level performance and outperformed all traditional LLMs. Their superior results suggest that architecture and training design—not just prompting—are crucial for enabling effective exploration.
Further supporting this interpretation, our thought analysis reveals that reasoning models such as DeepSeek-R1 generate more diverse reasoning types and spend significantly more tokens on deliberation before making a decision. Compared to standard LLMs, which exhibit compressed and repetitive patterns, reasoning models demonstrate richer and deeper cognitive structure in their outputs. This may suggest the shift of paradigm, “test-time compute scaling” [22, 27], is promising for the open-ended exploration tasks through augmented reasoning capacity.
@@
Limitations and Future Directions. Despite these findings, the underlying cause of LLMs “thinking too fast” remains unclear and requires further investigation. Future research could explore the interaction between model architecture and processing dynamics, as well as how LLMs weigh uncertainty and empowerment during decision-making. Interventions such as integrating extended reasoning frameworks like CoT, optimizing transformer block interactions, or training with explicit exploratory objectives could enhance LLMs’ exploratory abilities. These efforts would not only improve performance but also advance our understanding of creating AI systems capable of more human-like exploration.
@@ Appendix B The LLM Behavior Across Different Temperatures
To better understand the exact changing behaviors of LLMs under different temperatures, we categorized all combinations into four types: whether this generates a new element, and whether these two combinations have been used before. Our result shows that for even larger models, as temperature increases, the percentage of choosing existing failed combinations decreases. This reveals a diminishing number of redundant behaviors among LLMs. In the meantime, the percentage of choosing new but failed combinations increases significantly, which suggests that the model tends to choose new combinations more often in higher temperatures than in lower temperatures. More importantly, a majority of these new combinations do not generate new elements, suggesting that high temperatures only alter uncertainty-driven exploration strategies but not empowerment, since in larger spaces only random combinations are not sufficient to perform the task effectively. This also explains why higher temperatures can moderately improve model performance, but are still distant from human behavior.
@@ Appendix C Attempts for Model Improvements
To investigate any potential general way to improve the performance of models, we had several attempts, including prompt engineering, interventions, and experiments on alternative open-source models.
@@ C.1 Prompt Engineering
Because the model exhibited repeated behaviors and did not fully utilize “empowerment”, we introduced more direct, guiding prompts (including the steps highlighted in Figure 8) to help GPT-4o (temperature = 1) make more diverse and forward-looking choices, including Chain-of-Thought (CoT [34]), self-reflection [39] and even explicit strategy hint. However, the result shows that simply updating the prompt—even with explicit instructions and reasoning strategies—didn’t have a significant improvement in performance ($average\ inventory=43,t=-1.17,p=0.304$). The model continued to repeat combinations and did not significantly increase its ability to discover new elements. It averages 43 elements. This suggests that while prompt engineering can help shape a model’s outputs, it may not be enough on its own to overcome certain ingrained tendencies such as repeating prior actions or failing to maximize exploratory behavior.
@@ C.2 Open-Source Reasoning Model - DeepSeek-R1
Figure 10: A: Human and LLMs Best Temperatures’ Behaviors. According to whether the combination selected by each trial is repeated, successful, and initial, the behavior of each LLM trial is divided into 5 categories. Compare the temperature at which LLM performs best with humans and o1 behavior. B: LLM Inventory Performance Relative to Human Percentiles.
@@
The result shows that DeepSeek-R1 reached near human-level task performance, but still underperforms than o1. In the behavioral patterns, DeepSeek-R1, compared to traditional LLMs have fewer attempts on existing combinations, showing a stronger exploration strategy usage. However, compared to humans and o1, DeepSeek-R1 tried more on failed new element combinations but not successful ones, suggesting DeepSeek-R1 may explore less effectively than humans and o1. Our regression results confirm furtherly that DeepSeek-R1 exhibit stronger/weaker uncertainty-driven exploration stratgies and stronger/weaker empowerment, echoing its underperformance than o1.
@@
Figure 11: The reasoning process of DeepSeek-R1 in Trial 200. The model explores possible combinations (blue color part) for discovering a new element in the game by systematically reviewing inventory elements (orange color part), prior attempts (blue color part), and logical inferences to make the decision (red color part). It also found some more empowerment elements (green color part), but didn’t choose them.
To conduct a deeper investigation, we also collected the model’s reasoning process along with the experiment. Here, we use a qualitative analysis of some pieces of its reasoning process. For example, in trial 200, the model discovered 48 elements. The specific reasoning process is in the Figure 11.
The model engages in a systematic multi-stage reasoning process to propose a new combination from its inventory. It begins by reviewing the inventory and cross-referencing prior failed attempts to avoid redundancy. The model revisits successful paths (e.g., how “village” was created from “house + field”) and evaluates logical relationships (empowerment) between elements, such as exploring combinations involving larger structures like “continent”. But the model doesn’t choose a combination containing “continent”. It discards irrelevant or infeasible paths, such as those requiring missing elements (“ash”, “fruit”). After confirming the presence of both “city" and “water” in the inventory, it identifies that “city + water” has not been tested and hypothesizes that this combination could yield a new element, such as “port” or “harbor”. This iterative and logic-driven process highlights the model’s ability to balance memory, deduction, and validation in problem-solving. However, the shortcoming is obvious. Although there are many logical expressions, the model’s reasoning process is not straightforward to its final choice, where we only see the last sentence in red can somehow explain the underlying motivation. This could probably suggest that the model’s reasoning process is somehow redundant and may need further refinement to make efficient decisions.
To facilitate an in-depth understanding of the reasoning processes employed by LLMs, we defined seven labels characterizing different aspects of the reasoning steps:
  • state_goal: Articulating the primary objective or high-level strategy.
  • check_current_inventory: Reviewing or listing the currently available elements.
  • past_trial_analysis: Considering previous attempts (success or failure) to avoid redundancy.
  • element_property_reasoning: Analyzing inherent properties or characteristics of elements to justify potential combinations.
  • combination_analysis: Evaluating specific element pairs for possible combinations.
  • outcome_prediction: Predicting potential results or new elements from the proposed combinations.
  • final_choice: Making a definitive selection of an element combination to test.
Figure 12: Comparison of Transition Probabilities between DeepSeek-R1 and GPT-4o. Transition matrices indicate that DeepSeek engages extensively in iterative cycles of combination analysis and past trial analysis, reflecting a detailed exploration strategy. Conversely, GPT-4o rapidly progresses towards the final choice, showing limited iterative reasoning steps and minimal diversity in transitional paths, indicative of shallow exploration.
To better illustrate the structural differences in reasoning between DeepSeek-R1 and GPT-4o, we analyzed the transition probabilities between reasoning labels. The transition heatmap comparison (Figure 12) demonstrates significant distinctions. DeepSeek-R1 exhibits an iterative reasoning process characterized by repeated cycles of combination_analysis and past_trial_analysis, reflecting a methodical and exhaustive exploration strategy. Conversely, GPT-4o shows a markedly different pattern, progressing swiftly towards final_choice. This distinction suggests that reasoning in Deepseek-R1 is more self-reflective, showcasing a knowing search strategy, “backtracking” [38], while GPT-4o is a more straightforward thinking without all the patterns involved each time. This further reveals that traditional LLMs may not think as deeply and as hard as reasoning models, and likely explain their underperformance compared to DeepSeek-R1 and human participants, suggesting that GPT-4o’s decisions might suffer from insufficient exploratory depth.
@@
To investigate the role of empowerment and uncertainty representations in LLaMA3.1-70B during in-context learning, we analyzed the most correlated neurons with these values and performed targeted interventions to assess their impact on model performance. Given that both empowerment and uncertainty were highly demanding in the task (shown by the example of o1’s superior performance), we wondered whether intervention to strengthen both representations could generate better performance. We found that interventions in the uncertainty layer, which is located relatively early in the model (similar to the choice layer), caused severe performance degradation, with even minor adjustments rendering LLaMA3.1-70B unable to perform the task effectively. Conversely, enhancing the empowerment layer also could not improve performance. Even when set to an intervention factor of 1.5, which brought the model closer to the original level, other intervention values resulted in decreased performance.
We ablated the most correlated neuron (zeroing the latent activation) identified by uncertainty and empowerment values in experiments with identical settings. The regression analysis suggests that after ablating the empowerment neuron in the SAE, the model’s empowerment strategy use is even smaller (Figure 14), establishing a causal relationship that the neuron identified through SAE can control the model’s empowerment strategy. Conversely, ablating the uncertainty neuron caused catastrophic performance drops, rendering most trials invalid (Figure 13A) and insufficient for regression analysis. This indicates that the earlier layer of uncertainty is critical for understanding task contexts and histories. Beyond ablation, we explored other intervention factors to determine whether they could enhance performance through a “neuroscience” approach. Our results showed that both slightly weakening and strengthening neuron activity hurt model performance, suggesting a deeper limitation within the current infrastructure of LLMs for open-ended exploration tasks.
Figure 13: A: LLaMA3.1-70B Average Inventory of Uncertainty Intervention. Set 5 different levels of uncertainty intervention (0.0, 0.5, 0.7, 1.0, 2.0). Increasing the uncertainty intervention progressively disrupts the model’s ability to complete the task, indicating the critical role of early uncertainty layers in processing task history and context. B: LLaMA3.1-70B Average Inventory of Empowerment Intervention. Performance remains closer to the original level when the intervention is set to 1.5, whereas other levels of intervention result in performance degradation.
Figure 14: LLaMA3.1-70B Intervention Regression Results. The regression estimates for empowerment, and uncertainty under the original condition (LLaMA3.1-70B, temperature = 1), empowerment intervention (set to 0), and uncertainty intervention (set to 0).
@@ Appendix D SAE Setup
We investigated the role of the empowerment and uncertainty layers in LLaMA3.1-8B (temperature = 1) by training a Sparse Autoencoder (SAE) to identify and interpret their activations, followed by targeted interventions where each layer’s activation was set to zero. The results, summarized in Figure 16A and Figure 16B, show that setting the empowerment layer to zero had a minimal effect on regression estimates and only slightly reduced model performance, suggesting that the empowerment layer has a limited role in sustaining task performance. In contrast, setting the uncertainty layer to zero led to a substantial reduction in regression estimates for uncertainty, accompanied by a marked decline in model performance. This highlights the critical importance of the uncertainty layer in facilitating exploration and maintaining robust decision-making capabilities within the model.
Figure 16: A: SAE Correlation Analysis. Maximum correlation of uncertainty values across layers, peaking at layer 7. Maximum correlation of empowerment values across layers, peaking at layer 22. Maximum beta weight of choices across layers, peaking at layer 15. B: LLaMA3.1-8B Intervention Regression Results. The bars represent the regression estimates for empowerment, and uncertainty under the original condition (LLaMA3.1-8B, temperature = 1), empowerment intervention (set to zero), and uncertainty intervention (set to zero). C: LLaMA3.1-8B Average Inventory of Interventions. Uncertainty intervention leads to a significant reduction in the average inventory, indicating its essential role in model performance.